OpenStudy (anonymous):

6/15+2/15 reduce to lower terms

OpenStudy (anonymous):



hero (hero):

I think the instructions should say "combine" terms rather than "reduce"

OpenStudy (anonymous):

@hero it should be combined first and then it is reduced... it is a rule

OpenStudy (anonymous):

This can not be reduced further.

OpenStudy (anonymous):

got it. thanks

hero (hero):

After combining, there won't be anything to reduce.

OpenStudy (anonymous):

(Although, technically, "reduce" means to find an equivalent fraction and not necessarily a simpler one.)

OpenStudy (anonymous):

1/7+5/14=?

OpenStudy (anonymous):

lol that will be the answer.... and it will not be reduced

hero (hero):

Hint: Multiply the first fraction by 2/2

OpenStudy (anonymous):

1/7 + 5/14 = 2/14 + 5/14. (I first reduced 1/7 to 2/14 to get a common denominator).